OpenStudy (anonymous):

how do you find the area and perimeter of a square with a diagonal of 10 inches

OpenStudy (anonymous):

use properties of special right triangles, a square's diagonal cuts the the square into 2 isosceles right triangles

OpenStudy (anonymous):

so do i use 30-60-90 or 45-45-90?

OpenStudy (anonymous):

45-45-90

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

 using pythogaraus theorem 2x^2=100.Which means x is \[\sqrt{50}\] so now are is square of that which is 50 and perimeter is \[4\sqrt{50}\]